OpenStudy (anonymous):

Verify the identity cot(theta - pi/2)= -tan(theta) .

OpenStudy (anonymous):

Use these facts:\[ \tan x = \frac{\sin x}{\cos x} \\ \\ \cot x = \frac{\cos x}{\sin x} \\ \sin (x - \pi/2) = \sin (-( \pi/2 - x)) = -\sin (\pi/2 - x)=-\cos x \\ \cos( x - \pi/2) = \cos( -(x - \pi/2)) = \cos( \pi/2 - x ) = \sin x \]

OpenStudy (anonymous):

ok but I got stuck on how to start he problem then when I attempted to work it out it looks so wrong